OpenStudy (anonymous):

The formula for the amount with simple interest is A = P(1 + rt). Find t if A = 168, P = 150, and r = 0.08.

OpenStudy (king):

formula for simple interest is PNR/100

OpenStudy (king):

so, 168=150*0.08*t/100 t=16800/150*0.08 t=1400